the wavelength of an electron of energy 10KeV will be ?

Grade:12th pass

1 Answers

Parvin
6 years ago
Electrons follow de broglies eq.Lambda =h/mv Kinetic energy =1/2 mv^2KE=1/2×mv^2m(KE)=mv^2÷2×massM (KE)=m^2×v^2÷2=(p)^2÷2P=root (2M (ke))So lambda=h÷(root2m (KE))Putting the values6.63×10^-34÷root (18.2×10^-27)On solving...we get Wavelength =0.49×10^-47